OpenStudy (anonymous):

] Solve the equation x − 3 = 13 for x. Please show your work and solution.

OpenStudy (unklerhaukus):

\[x − 3 = 13\]add three to both sides \[x-3+3=13+3\]simplify\[\qquad\qquad=\]